OpenStudy (anonymous):

Please help me on a few of these. 1/6 = 64^4x-3 The generation time G for a particular bacteria is the time it takes for the population to double. The bacteria increase in population is shown by the formula G= t/3.3loga^p, where t is the time period of the population increase, a is the number of bacteria at the beginning of the time period, and P is the number of bacteria at the end of the time period. If the generation time for the bacteria is 2 hours, how long will it take 11 of these bacteria to multiply into a colony of 8764 bacteria?

OpenStudy (anonymous):

log(9x+5) = 0 solve

OpenStudy (anonymous):

log(x+9)-logx=3

OpenStudy (anonymous):

log4x+log7 =1

OpenStudy (anonymous):

log2x+log12=0

OpenStudy (anonymous):

you're using OS for a test? against the code of conduct

OpenStudy (anonymous):

its an online assignment that counts as a test grade

OpenStudy (anonymous):

than help me?

OpenStudy (anonymous):

key word HELP

OpenStudy (solomonzelman):

Which problem do you need help with right now ?

OpenStudy (anonymous):

the first one

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf log_{10}(9x+5) = 0 }\) when base is not specified it is 10, but in this case it doesn't matter, b/c \(\LARGE\color{blue}{ \bf log_{~anything~}(1) = 0 }\) what makes 9x+5=1 ?

OpenStudy (solomonzelman):

@danielle8008 , hello, are you there ?

OpenStudy (anonymous):

yes -4/9

OpenStudy (solomonzelman):

Yes, you are correct. Next you need help with number 2 ?

OpenStudy (anonymous):

yes please

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf log(x+9)-log(x)=3 }\) \(\LARGE\color{blue}{ \bf log(x+9)-log(x)=3~log~10 }\) \(\LARGE\color{blue}{ \bf log(x+9)-log(x)=log~10^3 }\) \(\LARGE\color{blue}{ \bf log(x+9)-log(x)=log~1000 }\) \(\LARGE\color{blue}{ \bf log(x+9)=log~1000 +log~x }\) \(\LARGE\color{blue}{ \bf log(x+9)=log~1000x }\) \(\LARGE\color{blue}{ \bf x+9=1000 x }\) ask anything if you don't get about what I did

OpenStudy (anonymous):

hmm ok the third one now

OpenStudy (solomonzelman):

\(\LARGE\color{red}{rule:~~~~ \bf B~log(A)=log(A^{B})}\)

OpenStudy (solomonzelman):

Ohh, i thought you didn't understand the third equation.... :)

OpenStudy (solomonzelman):

\(\LARGE\color{green}{ \bf log4x+log7 =1 }\) \(\LARGE\color{green}{ \bf log(4x \times 7 )=1 }\) \(\LARGE\color{green}{ \bf log(28x )=1 }\) \(\LARGE\color{green}{ \bf log(28x )=1~log10 }\) \(\LARGE\color{green}{ \bf log(28x )=log10^1 }\) \(\LARGE\color{green}{ \bf log(28x )=log10 }\) \(\LARGE\color{green}{ \bf 28x =10 }\)

OpenStudy (anonymous):

on of the answer choices is not 10/28 or 2.8

OpenStudy (anonymous):

*one of

OpenStudy (solomonzelman):

are you telling me that 10/28 =2.8 ? Do not confuse 10/28 with 28/10.

OpenStudy (anonymous):

one is 280

OpenStudy (solomonzelman):

you got to be kidding me right now.

OpenStudy (solomonzelman):

10/28=5/14 ≈0.36

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

duh sorry

OpenStudy (anonymous):

brain fart

OpenStudy (solomonzelman):

so do you still need help with this problem (number 3 ) ??

OpenStudy (anonymous):

i understand that one

OpenStudy (anonymous):

can you help me with two other ones i did not post?

OpenStudy (solomonzelman):

you actually have 1 left that you posted :)

OpenStudy (anonymous):

The number of bacteria present in a culture after t minutes is given as b=100e^kt There are 6639 bacteria present after 5 minutes. Find k. Explain how you solve this problem.

OpenStudy (solomonzelman):

Hold on, you didn't do \(\LARGE\color{green}{ \bf log2x+log12=0 }\)

OpenStudy (anonymous):

oh yeah

OpenStudy (solomonzelman):

\(\LARGE\color{green}{ \bf log2x+log12=0 }\) \(\LARGE\color{green}{ \bf log2x=-log12 }\) \(\LARGE\color{green}{ \bf log2x=(-1)~log12 }\) \(\LARGE\color{green}{ \bf log2x=log(12^{-1}) }\) \(\LARGE\color{green}{ \bf log2x=log(1/12) }\) \(\LARGE\color{green}{ \bf 2x=1/12 }\)

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

tell me the answer to this problem, before we go on.

OpenStudy (anonymous):

.357

OpenStudy (solomonzelman):

What about an exact answer ?

OpenStudy (anonymous):

.36 ?

OpenStudy (anonymous):

im not sure what you are asking

OpenStudy (solomonzelman):

I am not telling you to round up the decimal, I am asking for the exact x value, in terms of a fraction.

OpenStudy (anonymous):

1/12/2

OpenStudy (solomonzelman):

that is not a fraction... or not the type of fraction I want.

OpenStudy (anonymous):

.08/2

OpenStudy (anonymous):

can we go to the next one now? i chose .36 from the answer options on the assignment

OpenStudy (solomonzelman):

The number of bacteria present in a culture after t minutes is given as b=100e^kt There are 6639 bacteria present after 5 minutes. Find k. Explain how you solve this problem. \(\LARGE \color{blue}{ \rm b=100e^{kt} }\) \(\LARGE \color{blue}{ \rm 6639=100e^{~k\times 5} }\) \(\LARGE \color{blue}{ \rm 6639=100e^{~5k} }\) I got to go now. I am tired also. I am suggesting to plug in the approximate 2.7 for e. and solve using log.

OpenStudy (anonymous):

thanks